# Fine-tuning Whisper on Low-Resource Languages for Real-World Applications

### v2

@@ Abstract
This paper presents a new approach to fine-tuning OpenAI’s Whisper model for low-resource languages by introducing a novel data generation method that converts sentence-level data into a long-form corpus, using Swiss German as a case study. Non-sentence-level data, which could improve the performance of long-form audio, is difficult to obtain and often restricted by copyright laws. Our method bridges this gap by transforming more accessible sentence-level data into a format that preserves the model’s ability to handle long-form audio and perform segmentation without requiring non-sentence-level data. Our data generation process improves performance in several real-world applications and leads to the development of a new state-of-the-art speech-to-text (STT) model for Swiss German. We compare our model with a non-fine-tuned Whisper and previous state-of-the-art Swiss German STT models, where our new model achieves higher BLEU scores. Our results also indicate that the proposed method is adaptable to other low-resource languages, supported by written guidance and code that allows the creation of fine-tuned Whisper models, which keep segmentation capabilities and allow the transcription of longer audio files using only sentence-level data with high quality.
@@ 1 Introduction
Swiss German refers to the dialects spoken in the German-speaking regions of Switzerland. Due to the limited number of speakers, linguistic resources are scarce, also because Swiss German exists only as a spoken language, without any formal grammar or standardized written form. As a result, Swiss German STT systems are typically formulated as speech translation tasks, where Swiss German audio is transcribed into standard German text Plüss et al. (2021, 2022, 2023).
@@
Table 1: Comparison of a rapidly spoken Swiss-German saying Straub (2024), transcribed by Whisper Large-v2. The fine-tuned model fails to predict timestamps and performs worse than the original Large-v2.
Input Audio
Ich zeig der, wo de Bartli de Moscht holt.
Whisper Large-v2
[00:00:08] Ich zeige dir, wo Bartli den Most holt. [00:00:11]
Whisper Large-v2 (fine-tuned on sentences)
Ich zeige dir, wo es die Bartli in den Most holt.
@@ 2 Related work
Despite the challenge of scarce data resources, recent advancements in speech translation for Swiss German have been substantial, driven in part by recent collections of high-quality sentence-level datasets from Swiss parliaments minutes and crowdsourcing initiatives such as SwissDial Dogan-Schönberger et al. (2021), the Swiss Parliaments Corpus Plüss et al. (2021), SDS-200 Plüss et al. (2022), and STT4SG-350 Plüss et al. (2023).
@@ 3.1 Data Generation
In Table 5 we show the influence of these data generation strategies on different datasets. In Figure 2 the general approach is shown. The beginning and the end of speech are detected and then, when concatenating sentences together, we can either:
@@ 3.2 Training Details
Max Learning Rate $2.0\times 10^{-4}$
@@ 3.3 Train, Validation and Test Data
Based on the unabated rise of the curve in Figure 3, it’s reasonable to deduct that more hours of data could improve the model further.
@@ 4.1 Evaluation
In all experiments, the computed metrics are derived from the WhisperX [^4] https://github.com/m-bain/whisperX versions of the models using fp16 precision. The beam size is configured to 5, VAD is enabled and the language tag is DE. For the sentence-level datasets, we report the WER or the BLEU metric or both, if the layout allows it. For the long-form test dataset we replace the BLEU metric with SubER Wilken et al. (2022); this allows us to incorporate a metric that also assesses the segmentation quality. Before calculating the metrics, all sentences are transformed to lowercase, and punctuation is removed. For the BLEU metric we specifically use sacreBLEU Post (2018) with default parameters.
@@ 4.2 Base Results
Table 4: Performance of the original Whisper models on various test sets. SubER is shown on the long-form SRG data. *
@@ 4.3 Segmentation Forgetting & Out-of-Distribution Performance Degradation
Table 5: Comparing BLEU and SubER metrics for Whisper Large-v2 fine-tuned on various data augmentation methods versus fine-tuned on our generated long-form dataset. As seen in the SRG and Dataset-A results, the sentence-level model performs significantly worse than with long-form data.
@@
While the sentence-level model attains with 77.38 the highest BLEU score on the STT4SG-350 dataset, it performs poorly on longer audio sequences in SRG and Dataset-A, especially when the segmentation quality is taken into account. This underscores its limitations with out-of-distribution data. In comparison, models trained with the generated long-form dataset — especially ’All’ — demonstrate better generalization, maintaining higher BLEU scores and lower SubER scores across different datasets, nearly beating Whisper Large-v2 on the SRG dataset on BLEU.
The pre-trained Large-v2 model without fine-tuning, outperforms the sentence-level model on long-form datasets. This suggests that fine-tuning exclusively on sentence-level data processed of any kind into long-form, degrades the performance on unseen datasets. Incorporating long-form audio and datasets with diverse distribution into the training process is essential for preserving segmentation capabilities and ensuring robust performance across diverse data distributions.
@@ 4.4 Overcome Shortcomings
This leads to a model fine-tuned with our generated corpus based on the three sentence-level datasets (SPC, SDS-200, and STT4SG-350), the pseudo-labeled dataset SRG (PL) based on Swiss German TV shows, and the German part of the Common Voice 13 training data, concatenated as described above. This strategy significantly improves the model performance, as shown in Table 6, and leads to a new state-of-the-art model for Swiss German speech-to-text, referred to as SOTA.
Table 6: Our new SOTA model, fine-tuned with data generation method ’All’ and supplemented with SRG (PL) and Common Voice 13 de, compared to the original Whisper Large-v2
@@ 4.5 Dialect Comparison
The results in Table 7 show large differences in the performance of the original Large-v2 model across the different dialect regions. In contrast, our fine-tuned SOTA model exhibits improved WER over a much narrower range across all dialects. For the reader, the improvements over other Swiss-German ASR models Schraner et al. (2022) are also shown.
Table 7: WER for Swiss German dialects on the STT4SG-350 test set for selected models; XLS-R and TF are older models published in Schraner et al. (2022).

### v3

@@ Abstract
This paper presents a new approach to fine-tuning OpenAI’s Whisper model for low-resource languages by introducing a novel data generation method that converts sentence-level data into a long-form corpus, using Swiss German as a case study. Non-sentence-level data, which could improve the performance of long-form audio, is difficult to obtain and often restricted by copyright laws. Our method bridges this gap by transforming more accessible sentence-level data into a format that preserves the model’s ability to handle long-form audio and perform segmentation (by predicting timestamps) without requiring non-sentence-level data. Our data generation process improves performance in several real-world applications and leads to the development of a new state-of-the-art speech-to-text (STT) model for Swiss German. We compare our model with a non-fine-tuned Whisper and previous state-of-the-art Swiss German STT models, where our new model achieves higher BLEU scores. Our results also indicate that the proposed method is adaptable to other low-resource languages, supported by written guidance and code that allows the creation of fine-tuned Whisper models, which keep segmentation capabilities and allow the transcription of longer audio files using only sentence-level data with high quality.
@@ 1 Introduction
Swiss German refers to the dialects spoken in the German-speaking regions of Switzerland. Due to the limited number of speakers, linguistic resources are scarce, also because Swiss German exists only as a spoken language, without any formal grammar or standardized written form. As a result, Swiss German STT systems are typically formulated as speech translation tasks, where Swiss German audio is transcribed into standard German text Plüss et al. (2021, 2022, 2023). Generally, "Swiss German transcription" or "Swiss German ASR" refers to converting spoken Swiss German directly into written Standard German, combining transcription and translation in a single step.
@@
Table 1: Comparison of a rapidly spoken Swiss-German saying Straub (2024), transcribed by Whisper Large-v2. The model, fine-tuned on sentence-level data, fails to predict timestamps and performs worse than the original Large-v2.
Input Audio
Ich zeig der, wo de Bartli de Moscht holt.
Whisper Large-v2
[00:00:08] Ich zeige dir, wo Bartli den Most holt. [00:00:11]
Whisper Large-v2 (fine-tuned on sentences)
Ich zeige dir, wo es die Bartli in den Most holt.
@@ 2 Related work
Despite the challenge of scarce data resources, recent advancements in speech translation for Swiss German have been substantial, driven in part by recent collections of high-quality sentence-level datasets from Swiss parliaments minutes and crowdsourcing initiatives such as SwissDial Dogan-Schönberger et al. (2021), the Swiss Parliaments Corpus SPC Plüss et al. (2021), SDS-200 Plüss et al. (2022), and STT4SG-350 Plüss et al. (2023).
@@ 3.1 Data Generation
Later in the study, in Table 5, we show the influence of these data generation strategies on different datasets. In Figure 2 the general approach is shown. The beginning and the end of speech are detected and then, when concatenating sentences together, we can either:
@@ 3.2 Training Details
Max. Learning Rate $2.0\times 10^{-4}$
@@ 3.3 Train, Validation and Test Data
Based on the unabated rise of the curve in Figure 3, it’s reasonable to deduct that more training data will improve model performance further.
@@ 4.1 Evaluation
In all experiments, the computed metrics are derived from the WhisperX [^4] https://github.com/m-bain/whisperX versions of the models using fp16 precision. The beam size is configured to 5, VAD is enabled and the language tag is DE. For the sentence-level datasets, we report the WER or the BLEU metric or both, if the layout allows it. For the long-form test dataset we replace the BLEU metric with SubER Wilken et al. (2022); this allows us to incorporate a metric that also assesses the quality of segmentation by timestamp prediction. Before calculating the metrics, all sentences are transformed to lowercase, and punctuation is removed. For the BLEU metric we specifically use sacreBLEU Post (2018) with default parameters.
@@ 4.2 Base Results
Table 4: Performance of the original Whisper models on various sentence-level test sets. SubER is shown on the long-form SRG data.
@@ 4.3 Segmentation Forgetting & Out-of-Distribution Performance Degradation
Table 5: Comparison of BLEU and SubER metrics for Whisper Large-v2, fine-tuned on sentence-level (first row) and differently generated long-form data (Large-v2 refers to the original model without fine-tuning). The results on SRG and Dataset-A show that the sentence-level fine-tuning performs noticeably worse compared to training on long-form data.
@@
While the sentence-level model attains with 77.38 the highest BLEU score on the STT4SG-350 dataset, it performs poorly on longer audio sequences in SRG and Dataset-A, especially when the segmentation quality is taken into account. This underscores its limitations with out-of-distribution data. In comparison, models trained with the generated long-form dataset — especially ’All’ — demonstrate better generalization, maintaining higher BLEU scores and lower SubER scores across different datasets, nearly reaching the original Whisper Large-v2 on the Dataset-A on BLEU.
The original Large-v2 model without fine-tuning, outperforms the sentence-level model on long-form datasets. This suggests that fine-tuning exclusively on sentence-level, which was data processed into long-form, degrades the performance on unseen datasets. Incorporating long-form audio and datasets with diverse distribution into the training process is essential for preserving segmentation capabilities and ensuring robust performance across diverse data distributions.
@@ 4.4 Overcome Shortcomings
This leads to a model fine-tuned on long-form audio with our generated corpus based on the three sentence-level datasets (SPC, SDS-200, and STT4SG-350), the pseudo-labeled dataset SRG (PL) based on Swiss German TV shows, and the German part of the Common Voice 13 training data, concatenated as described above. This strategy significantly improves the model performance, as shown in Table 6, and leads to a new state-of-the-art model for Swiss German speech-to-text, referred to as SOTA.
Table 6: Our new Whisper Large-v2 based SOTA model, fine-tuned on long-form audio created with data generation method ’All’ and supplemented with SRG (PL) and Common Voice 13 de, compared to the original Whisper Large-v2
@@ 4.5 Dialect Comparison
The results in Table 7 show large differences in the performance of the original Large-v2 model across the different dialect regions. In contrast, our fine-tuned SOTA model exhibits improved WER over a much narrower range across all dialects. For the reader, the improvements over other Swiss-German ASR models based on Wav2Vec (XLS-R) and Transformer (TF) architectures Schraner et al. (2022) are also shown.
Table 7: WER for Swiss German dialects on the STT4SG-350 test set for selected models; XLS-R and TF are older models based on Wav2Vec (XLS-R) and Transformer (TF) architectures Schraner et al. (2022).